1. Using a 0.20 M phosphate buffer with a pH of 6.7, you add 0.79 mL of 0.53 M HCl to 40. mL of the buffer. What is the new pH of the solution? (Enter your answer to three significant figures.)

2. Using a 0.20 M phosphate buffer with a pH of 6.7, you add 0.79 mL of 0.53 M NaOH to 40. mL of the buffer. What is the new pH of the solution? (Enter your answer to three significant figures.)
Answer #1

Solution - Given that: 0 phosphate beffer = = 0-90 m. 0.90 M. . Hel - Ho.ml Giveni Hendron Cuention :- . pilt apta + log ((A-[A] = [HA] Toral Concentration of Polfer = 0•JO M. (+14] + [HA] 20:30 M. Acetic Concentration, (WA) = 0.20 0.0943 M 2112 AcceHendoyon ewation pH = 4.75 + log 0.04901 ) O 055785 -0.054425TH1896 4.75 & H.69 Ho7 =4075 -0.05 430004 tot CH3400+ H2O. fctzRemaining mole of CH3.com = 0.003772 -0.00004187 0.0037301 initial mole of citycool4 = 0.1057 mol X40 loco mi 0.0001057 X40 TTotal volume = 75.79 ml CH2COOH 0.05591 = 0.004238 : 71000 75.79 (CH2600] -0.0036136 xas 0.050320 X 75.79 1000 Henderson Caua